OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. Through (1, −2); parallel to the line x + 2y = 6

OpenStudy (loser66):

@anikhalder

OpenStudy (anonymous):

@anikhalder you got this

OpenStudy (anikhalder):

first let's find the slope of the given line..let's first convert it to y=mx + c form...m is the slope of the given line..let's first find that out

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

I like that idea

OpenStudy (loser66):

go ahead @anikhalder

OpenStudy (anikhalder):

what will be the equation in y=mx+b form @Stacy13

OpenStudy (anonymous):

got it

OpenStudy (anikhalder):

cool..but if we have to find a line perpendicular then we have to just interchange the coefficients of x and y with a -ve sign or something like that i presume?...for the trick but the long method is cool to get correct answer @Loser66

OpenStudy (loser66):

@anikhalder to perpendicular line, we can turn it to // by convert the slope to -1/m, and then apply the trick, why not? right?

OpenStudy (anikhalder):

yes..we can definitely do that...that's the lengthier process but suppose we are given an eqn 3x+4y=6...we have to find a line perp to this: then cant we do it something like 4x-3y=c ?

OpenStudy (loser66):

hehehe, anyway, I am cheating, sorry for that. You are 100% right, perfectly correct. Students shouldn't cheat. hehehe. My bad.

OpenStudy (anikhalder):

@Loser66 ..please tell me whether this method works for perp lines..the one i said..i'm not sure

OpenStudy (anikhalder):

okay...thanks

OpenStudy (anikhalder):

that's not cheating...it's just that you are applying a general form that you have valid proof to account for..it's like using the formula to solve a quad eqn instead of factorizing it...but i think as you said...we should only apply the trick that you used when we are pretty sure of the longer method

OpenStudy (loser66):

thnks

OpenStudy (anikhalder):

thanks to you....and not to me..:)